bill533:

A circle is centered at C(−1,−3) and has a radius of 6.Where does the point P(-6,-6) lie? A. Inside the circle B. on the circle C. outside the circle

bill533:

@563blackghost

bill533:

A. correct ?

563blackghost:

Yup

bill533:

Given points A(-2, -4), and B (4, 8), which of the following represents the midpoint of segment AB? A. (-3, -2) B. (2, 4) C. (-6, -4) D. (1, 2)

563blackghost:

What answer do you think et is?

bill533:

i think D if i'm not mistaken

563blackghost:

Correct cx

bill533:

it is ?

563blackghost:

yes

bill533:

Given points A(1, 3) and B (3, 7), which of the following represents the distance of AB¯? A.√20 B. √116 C. √14 D. √6

563blackghost:

how bout dis one? what do you think the answer is?

bill533:

c.

563blackghost:

Mmm not quite. Use distance formula. \(\large\bf{d=\sqrt{(7-3)^{2}+(3-1)^{2}}}\)

bill533:

d = 2√5

563blackghost:

well yes, but the answers are simplified down to radical form, so it would be just \(\sqrt{20}\).

563blackghost:

are not*

bill533:

oh ok

bill533:

A.

563blackghost:

yes

bill533:

The equation of the line that has a slope of -1 and goes through (-3,0) is: A. y = -x B. y = -x - 3 C. y = x - 3 D. y = -x + 3

563blackghost:

what answer do you think this one is?

bill533:

B

563blackghost:

correct

bill533:

1 attachment
bill533:

A. In step 1, she compares the distances between the wrong pairs of points. B. In step 2, she incorrectly calculates the distance between (3, 1) and(3, –4) . C. In step 5, she simplifies incorrectly. D. In step 4, she makes a sign error.

bill533:

i think the answer is D

563blackghost:

I would agree.

bill533:

that was incorrect

563blackghost:

welp, which was the answer?

bill533:

It didn't say

563blackghost:

oh i see it was A

563blackghost:

I didnt read it correctly, instead of comparing the points that she is trying to prove with the center, she compared it with the other point on the graph. Sorry.

bill533:

oh well What is the equation of the line that is perpendicular to this line and passes through (-6, -2)? https://learning.k12.com/content/enforced/437405-COF_ID125380/G-GPE.B.5%20Q2.JPG?_&d2lSessionVal=gIqbusDzQX5eWYCxepgrlqzgn A. y = −2/3x − 6 B. y = −2/3x − 2 C. y = 3/2x +7 D. y = 3/2x − 4

563blackghost:

Which do you think is the answer to dis one?

bill533:

D

bill533:

or C

563blackghost:

Mmm which are you more confident in?

bill533:

D

563blackghost:

Not quiet. The slope is correct but the y-intercept is not. You need to plug in your point in \(\bf{y=\frac{3}{2}x+b}\) to find `b`. \(\large\bf{-2=\frac{3}{2}(-6)+b}\) What does b equal?

bill533:

b = 7

563blackghost:

Correct, so it is C.

bill533:

kk let me start a new question